OpenStudy (rock_mit182):

solve the compound inequality 3x + 6 < -3 or 5 – x ≤ 1

OpenStudy (rock_mit182):

\[3x +6-6 = 6+3 \]

OpenStudy (anonymous):

next comes 3x + 0 = 6 + 3

OpenStudy (rock_mit182):

\[-5+5-x \le 1+5\]

OpenStudy (anonymous):

then comes 3x = 9

OpenStudy (rock_mit182):

correcting the first one,\[3x< -9\]

OpenStudy (rock_mit182):

x< -3 or -x <=6

OpenStudy (rock_mit182):

is that right ?

OpenStudy (rock_mit182):

@zepdrix

OpenStudy (rock_mit182):

@wio

OpenStudy (anonymous):

second one is wrong

OpenStudy (rock_mit182):

Describe in words what the solution sets mean, and then display a simple line graph for each solution set

OpenStudy (anonymous):

Solution sets? It's the set of all numbers which are compatible solutions.

OpenStudy (rock_mit182):

\[-x \le 6 \]

OpenStudy (anonymous):

That is incorrect, try again.

OpenStudy (rock_mit182):

\[-x*(-1)\le (-1)*6\]

OpenStudy (rock_mit182):

\[x \ge -6\]

OpenStudy (rock_mit182):

what is the property of the last step ?

OpenStudy (anonymous):

No, first of all, where is the 6 from?

OpenStudy (rock_mit182):

\[5-x<1\] \[-5+5-x \le 1-5\]

OpenStudy (rock_mit182):

\[-x \le 4\]

OpenStudy (rock_mit182):

\[(-1)*-x \le 4*(-1)\]

OpenStudy (rock_mit182):

x>=-4 now what do i have to do ?

OpenStudy (anonymous):

that is correct

OpenStudy (rock_mit182):

is this a union compound inequality ?

OpenStudy (rock_mit182):

help me out with the writting part please !

OpenStudy (anonymous):

Or makes it sound like a union.

OpenStudy (anonymous):

If it is a union, then all numbers are solutions, it if is an intersection, then it ie between -4 and -3

OpenStudy (rock_mit182):

oh thanks a lot, buddy

OpenStudy (rock_mit182):

\[x<-3\] or \[x \ge -4\]

OpenStudy (rock_mit182):